The pole of 3y + 5x  +17=0 wrt to circle x^2+y^2+4x+6y+9=0 is ???
with solution

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Equations of polar is T = 0 
 
hence
 
let pole = (a,b)
 
ax + by + 2 (a+x) + 3(b+y) + 9 = 0
 
(a+2) x + (b+3) y + 2a + 3b +9 =0
 
on comparing
 
(a+2) /3 = (b+3)/5 = (2a+3b +9)/17
 
now solve for a and b
which is required.